OpenStudy (anonymous):

The diagonal of a square is 8 square root 2 units. Which is the perimeter of the square? 8 16 24 32

jimthompson5910 (jim_thompson5910):

The relationship between the diagonal of the square (d) and the side of the square (s) is given by \[d=s\sqrt{2}\] Solve for 's' to get \[s=\frac{d}{\sqrt{2}}\] Now plug in d=8 square root of 2 and simplify to find s \[s=\frac{d}{\sqrt{2}}\] \[s=\frac{8\sqrt{2}}{\sqrt{2}}\] \[s=8\] So the length of the side of the square is 8 units So the perimeter is 8*4 = 32 units.

OpenStudy (anonymous):

Thkank you (:

OpenStudy (anonymous):

Alternatively, Pythagorus for the square with side a 8 sqrt 2 = sqrt (2a^2) -> a = 8-> 32.

OpenStudy (anonymous):

A trig solution:\[\text{perimeter}=4*8\sqrt{2}\text{Sin}[45{}^{\circ}] \]

OpenStudy (anonymous):

Sorry.\[\text{perimeter}=4*8\sqrt{2}\text{Sin}[45{}^{\circ}]=32 \]